In America, household electrical outlets produce 110V of potential. If a lightbulb is designed to consume 110W of power, what is its resistance?
Ω

If that same lightbulb were plugged into a socket in europe, where household electrical outlets produce 220V of potential, how much power would it consume?
W
Answer #1

Solution 1)

We know, Power, P= V^2/ R

110 W=(110 V)^2/R

So, Resistance, R=110 ohms

=======

Solution 2)

Here, R=110 ohms

Voltage, V= 220 V

So, P= V^2/R= 220^2/110 = 440 Watrs

=======